On an average, a bus stops at a certain point every 15 minutes. Whats the probability that no bus will turn up in a single 15 minute interval?


1
Expert's answer
2022-05-31T11:03:43-0400

Let "X=" the number of buses: "X\sim Po(\lambda)."

Given "\lambda=1".


"P(X=0)=\dfrac{e^{-1}(1)^0}{0!}=\dfrac{1}{e}\approx0.36788"